# Scalar-induced gravitational waves including isocurvature perturbations with lattice simulations

### v1

@@ Abstract
Scalar-induced gravitational waves (SIGWs) open a unique window into early-universe physics. While their generation from adiabatic perturbations has been extensively studied, the contribution from isocurvature perturbations remains poorly understood. In this work, we develop a lattice simulation framework to compute the stochastic gravitational wave background from both pure isocurvature and mixed initial conditions. Our numerical results show excellent agreement with semi-analytical predictions in the pure isocurvature case. We further analyze multi-peak structures under general initial conditions and find that they closely match those produced in purely adiabatic scenarios. Additionally, we examine SIGWs in early matter-dominated eras, revealing that the peak amplitude and spectral slope are sensitive to the microphysical properties of the dominant field, such as the primordial black hole mass, abundance, or soliton decay rate. This study establishes lattice simulations as a robust tool for predicting SIGW spectra from complex primordial perturbations, with important implications for interpreting current and future gravitational wave observations.
@@ 1 Introduction
The properties of SIGWs are intimately tied to the nature of the primordial perturbations. While the majority of studies have focused on adiabatic perturbations [22, 23, 24, 25, 26], consistent with the models of single field inflation and strongly constrained by CMB observations [27, 28, 29], another fundamental class, isocurvature perturbations [30, 31, 32, 33, 34], remains a compelling and less explored possibility (see Ref. [35] for a review). Isocurvature modes, characterizing relative density contrasts between different cosmological components (e.g., radiation and dark matter), are a generic prediction of many multi-field inflation models [36, 37, 38, 39, 40], phase transition [41] and could be associated with the formation of primordial black holes (PBHs) [42, 43, 44, 45, 46, 47] or other exotic non-topological relics like Q-balls [48, 49]. Crucially, isocurvature perturbations, particularly those of Poissonian origin such as from PBHs, are inherently large, reaching amplitudes of $O(1)$ [43, 44]. Consequently, they can generate a substantial SIGW signal and must be carefully accounted for.
The exact semi-analytical formalism for isocurvature-induced SIGWs was first given in Ref. [50], which is strictly valid only under the condition that the relevant perturbation modes enter the horizon long before matter-radiation equality, while the complexity escalates significantly for mixed initial conditions (with both adiabatic and isocurvature perturbations). In contrast, lattice simulations offer a robust alternative, with an intrinsic advantage in accurately capturing effects from non-Gaussianity. While this method has previously been applied exclusively to adiabatic fluctuations [51], we systematically extend it in this work to incorporate isocurvature perturbations and mixed initial conditions.
@@ 2.1 Equations of motion
Given our focus on the early matter-dominated era (eMD), we accordingly assume that $Q^{\mu}$ originates entirely from the matter sector, and define it as
@@
$S_{ij}=4\partial_{i}\Phi\partial_{j}\Phi+2a^{2}\left(\rho_{(m)}\partial_{i}v_{(m)}\partial_{j}v_{(m)}+\frac{4}{3}\rho_{(r)}\partial_{i}v_{(r)}\partial_{j}v_{(r)}\right),$ (2.20)
and $T_{ij}^{~~lm}$ is the transverse-traceless projection operator defined by
@@
In our simulation, the scale factor $a(\tau)$ and conformal Hubble parameter $\mathcal{H}$ are evolved using Eq. (2.5), while Eq. (2.6) determines $\mathcal{H}^{\prime}$. The energy densities $\rho_{(m)}$ and $\rho_{(r)}$ are advanced via Eqs. (2.10) and (2.11), respectively. For the perturbative sector, Eqs. (2.14), (2.15), (2.16), (2.17), (2.18), and (2.19) govern the evolution of $\Phi,\delta\rho_{(m)},\delta\rho_{(r)},v_{(m)},v_{(r)}$ and $h_{ij}$, respectively. Simulations are performed in a comoving cubic box with periodic boundary conditions, following Ref. [51]. Spatial derivatives are computed using a Fourier pseudospectral method, and time integration is handled by a fourth-order Runge-Kutta scheme.
@@ 2.2 The initial conditions
$S_{k}(\xi)$ $=S_{i}(k),$ (2.28)
$\Phi_{k}(\xi)$ $=\Phi_{i}(k)\left(\frac{8}{5\xi^{3}}(\sqrt{1+\xi}-1)-\frac{4}{5\xi^{2}}+\frac{1}{5\xi}+\frac{9}{10}\right)$
$+S_{i}(k)\left(\frac{16}{5\xi^{3}}(1-\sqrt{1+\xi})+\frac{8}{5\xi^{2}}-\frac{2}{5\xi}+\frac{1}{5}\right),$ (2.29)
@@
$S_{k}^{\prime}(\xi)$ $=0,$ (2.30)
$\Phi_{k}^{\prime}(\xi)$ $=f(\tau)\Phi_{i}(k)\left(\frac{8}{5\xi^{3}}-\frac{1}{5\xi^{2}}+\frac{4}{5\xi^{3}\sqrt{1+\xi}}-\frac{24(\sqrt{1+\xi}-1)}{5\xi^{4}}\right)$
$+f(\tau)S_{i}(k)\left(-\frac{16}{5\xi^{3}}+\frac{2}{5\xi^{2}}-\frac{8}{5\xi^{3}\sqrt{1+\xi}}-\frac{48(1-\sqrt{1+\xi})}{5\xi^{4}}\right),$ (2.31)
@@
After defining the dimensionless power spectrum of Gaussian random fields $\Phi_{g}$ and $S_{g}$ [^1] Generalizing to non-Gaussian fields are straightforward [51].
@@
one can obtain the initial value of $\Phi_{i}$ and $S_{i}$ [51], where the equation of state parameter $w$ is defined as $\omega=p_{(r)}/(\rho_{(r)}+\rho_{(m)})$ and $\zeta_{g}$ is the gauge invariant curvature perturbation. For the initial relative velocity $V_{\mathrm{rel}}$, combining Eqs. (2.24), (2.25), and the exact solutions (2.28) and (2.29) shows that during radiation-domination or matter-domination,
@@ 3.2 The peak structures of general initial conditions
The number of such peaks is governed by momentum conservation, as represented by the step function in Eq. (3.2). However, in the isocurvature case, a suppression factor of the form $(k_{\mathrm{eq}}/k_{i})^{4}$ or $(k_{\mathrm{eq}}/k_{i})^{2}(k_{\mathrm{eq}}/k_{j})^{2}$ arises. This implies that certain peaks in the mixed initial condition exhibit an explicit dependence on $k_{\mathrm{eq}}$.
@@
To be specific, for the left panel of Fig. 2, we adopt $B_{1}=B_{2}=100$ in Eq. (3.4) to compute representative examples in isocurvature case; for the right panel, we choose $A=0.01$ in Eq. (3.2) for adiabatic perturbation and $B=100$ in Eq. (3.3) for isocurvature perturbation for numerical evaluation in mixed initial condition case.
@@ 4.1 SIGW from eMD isocurvature
where $\gamma\sim 0.2$ in the radiation-dominated (RD) era, $H_{f}$ is the Hubble parameter at the time of PBH formation, and $\beta$ denotes the fraction of PBHs at formation. Assuming that during the early radiation-dominated era (eRD), the energy density of PBHs is negligible compared to radiation at the time of PBH formation, and that initial density fluctuations on scales larger than $k_{\mathrm{UV}}^{-1}$ can be neglected, which means $\delta\rho_{\mathrm{PBH},f}+\delta\rho_{r,f}=0$. [^2] Initial density fluctuations arise when adiabatic perturbations are considered. In contrast, isocurvature perturbations emerge following PBH formation—typically after the peak-like adiabatic perturbation has entered the horizon. Therefore, within linear perturbation theory, the adiabatic and isocurvature modes could be treated independently. Therefore, the initial isocurvature perturbation is given by
@@
Figure 3: Left: The GWs energy spectra sourced by PBH isocurvature perturbations. We have set $k_{\mathrm{UV}}=40$ in our simulations. Right: The energy density fraction $\Omega_{i}=\rho_{i=m,r}/(\rho_{m}+\rho_{r})$ of PBH and radiation. The dashed lines denote $\Omega_{m}$, while the solid line describes $\Omega_{r}$. $t$ is the cosmic time.
As illustrative examples, we consider PBH masses of $M_{\mathrm{PBH}}=10^{-4}g$ and $2\times 10^{-4}g$, and initial abundances $\beta=5\times 10^{-2},10^{-2},5\times 10^{-3}$ in our simulations [^4] For larger PBH masses or smaller initial abundances, the values of $t_{\mathrm{eva}}$ or $\tau_{*}$ in the simulation become very large. Simulating such cases is very challenging, which may require switching to a logarithmic time evolution scheme for the equations., as shown in Fig. 3. It can be seen that a larger initial abundance or a larger PBH mass leads to a steeper power-law behavior near the peak, as illustrated in Fig. 4. This behavior may originate from the different decay times of PBHs and the duration of the PBH-dominated era, an interpretation that warrants further analytical and numerical investigation in future work.
Figure 4: Left: The GWs energy spectra sourced by PBH isocurvature perturbations for different PBH masses. Right: The GWs energy spectra sourced by PBH isocurvature perturbations for different PBH abundances. In both figures, we have multiplied a constant to make their ultraviolet part the same.
Figure 5: Left: The GWs energy spectra sourced by Q-balls isocurvature perturbations. We have set $k_{\mathrm{UV}}=40$ in our simulations. Right: The energy density fraction $\Omega_{i}=\rho_{i=m,r}/(\rho_{m}+\rho_{r})$ of Q-balls and radiation. The dashed lines denote $\Omega_{m}$, while the solid line describes $\Omega_{r}$. $t$ is the cosmic time. The initial abundance is set as $\beta=5\times 10^{-2}$.
@@ 4.2 SIGW from general initial conditions
A full treatment of general initial conditions is considerably more complex; here, we provide only a brief extension of the PBH-dominated case. In reality, isocurvature perturbations only emerge subsequent to PBH formation, typically well after the peak of the adiabatic perturbation has entered the horizon. Thus, one ought to initialize the adiabatic and isocurvature fluctuations at distinct times, a scenario beyond the scope of this work. Here, as an illustrative example, we simply superimpose the initial adiabatic and isocurvature contributions, while deferring the treatment of more realistic cases to future work. Assuming that PBHs are formed from a peak-like adiabatic perturbation centered at $k_{*}$, the ultraviolet cutoff scale $k_{\mathrm{UV}}$ can be related to $k_{*}$ via the identification $k_{*}=\mathcal{H}_{f}$
@@
For $\beta=10^{-2}$ and $\gamma=0.2$, one will obtain $k_{*}\simeq 2.7k_{\mathrm{UV}}$, which is shown in Fig. 6. As shown, the gravitational waves are initially generated by the adiabatic perturbation and are subsequently sourced by the isocurvature perturbation. The middle peak corresponds to the peak structure discussed in Subsection 3.2.
Figure 6: The GW energy spectrum sourced by the mixed initial condition at different time steps. The power spectrum of the initial adiabatic perturbation is the Gaussian bump power spectrum Eq. (3.2) with $A=0.01,\Delta=1/30$ and $k_{*}=2.7k_{\mathrm{UV}}=54$, while the initial isocurvature power spectrum is Eq. (4.5) with $k_{\mathrm{UV}}=20$. $\tau_{i}$ is the initial comoving time.

### v2

@@ Abstract
Scalar-induced gravitational waves (SIGWs) open a unique window into early-universe physics. While their generation from adiabatic perturbations has been extensively studied, the contribution from isocurvature perturbations remains largely unexplored. In this work, we develop a lattice simulation framework to compute the stochastic gravitational wave background from both pure isocurvature and mixed initial conditions. Our numerical results show excellent agreement with semi-analytical predictions in the pure isocurvature case. We further analyze multi-peak structures under general initial conditions and find that they closely match those produced in purely adiabatic scenarios. Additionally, we examine SIGWs in early matter-dominated eras, revealing that the peak amplitude and spectral slope are sensitive to the microphysical properties of the dominant field, such as the primordial black hole mass, abundance, or soliton decay rate. This study establishes lattice simulations as a robust tool for predicting SIGW spectra from complex primordial perturbations, with important implications for interpreting current and future gravitational wave observations.
@@ 1 Introduction
The properties of SIGWs are intimately tied to the nature of the primordial perturbations. While the majority of studies have focused on adiabatic perturbations [17, 75, 1, 41, 19], consistent with the models of single field inflation and strongly constrained by CMB observations [4, 12, 3], another fundamental class, isocurvature perturbations [44, 45, 46, 49, 14], remains a compelling and less explored possibility (see Ref. [30] for a review, and see Refs. [78, 54, 79] for some recent works). Isocurvature modes, characterizing relative density contrasts between different cosmological components (e.g., radiation and dark matter), are a generic prediction of many multi-field inflation models [22, 21, 56, 69, 51], phase transition [15] and could be associated with the formation of primordial black holes (PBHs) [37, 64, 24, 26, 27, 10, 65, 73] or other exotic non-topological relics like Q-balls [53, 52]. Crucially, isocurvature perturbations, particularly those of Poissonian origin such as from PBHs, are inherently large, reaching amplitudes of $O(1)$ [64, 24]. Consequently, they can generate a substantial SIGW signal and must be carefully accounted for.
The exact semi-analytical formalism for isocurvature-induced SIGWs was first given in Ref. [25], which is strictly valid only under the condition that the relevant perturbation modes enter the horizon long before matter-radiation equality, while the complexity escalates significantly for mixed initial conditions (with both adiabatic and isocurvature perturbations). In contrast, lattice simulations offer a robust alternative, with an intrinsic advantage in accurately capturing effects from non-Gaussianity. While this method has previously been applied exclusively to adiabatic fluctuations [81], we systematically extend it in this work to incorporate isocurvature perturbations and mixed initial conditions. We note in Ref. [32], the authors have applied a hybrid N-body and lattice simulation to study the GWs from structure formation in the early matter-dominated era, while our method has treated the matter as a fluid field, since the effects of structure formation are not our focus.
@@ 2.1 Equations of motion
Given our focus on the early matter-dominated era (eMD), we accordingly assume that $Q^{\mu}$ originates entirely from the matter sector, then, $Q^{\mu}$ can only be proportional to the four-velocity of the matter, thus, we can define it as
@@
For the PBH-dominated case, $Q$ should be proportional to the number denisty of PBHs $n_{\mathrm{PBH}}=\rho_{\mathrm{PBH}}/M_{\mathrm{PBH}}$ and their mass decay rate, which leads to
$Q=-n_{\mathrm{PBH}}u_{(m)}^{\mu}\partial_{\mu}M_{\mathrm{PBH}}.$ (2.10)
@@
$S_{ij}=$ $4\partial_{i}\Phi\partial_{j}\Phi+2a^{2}\left(\rho_{(m)}\partial_{i}v_{(m)}\partial_{j}v_{(m)}+\frac{4}{3}\rho_{(r)}\partial_{i}v_{(r)}\partial_{j}v_{(r)}\right)$
$=$ $4\partial_{i}\Phi\partial_{j}\Phi+6c_{s}^{2}\frac{\rho}{\rho_{r}}\partial_{i}\left(\frac{\Phi^{\prime}}{\mathcal{H}}+\Phi\right)\partial_{j}\left(\frac{\Phi^{\prime}}{\mathcal{H}}+\Phi\right)+6a^{2}c_{s}^{2}\rho_{m}\partial_{i}V_{\mathrm{rel}}\partial_{j}V_{\mathrm{rel}},$ (2.21)
where $V_{\mathrm{rel}}\equiv v_{(m)}-v_{(r)}$ and $T_{ij}^{~~lm}$ is the transverse-traceless projection operator defined by
@@
It is important to note that the term involving $V_{\mathrm{rel}}$ has been neglected in many of the previous literature, even though it can have a non-negligible effect in certain scenarios, as highlighted in Ref. [48]. In this work, we have explicitly tested its impact in Appendix B and find that it remains negligible in the PBH-dominated scenario discussed in Section 4. This result is consistent with the analysis presented in Ref. [24]. In our simulation, the scale factor $a(\tau)$ and conformal Hubble parameter $\mathcal{H}$ are evolved using Eq. (2.5), while Eq. (2.6) determines $\mathcal{H}^{\prime}$. The energy densities $\rho_{(m)}$ and $\rho_{(r)}$ are advanced via Eqs. (2.11) and (2.12), respectively. For the perturbative sector, Eqs. (2.15), (2.16), (2.17), (2.18), (2.19), and (2.20) govern the evolution of $\Phi,\delta\rho_{(m)},\delta\rho_{(r)},v_{(m)},v_{(r)}$ and $h_{ij}$, respectively. Simulations are performed in a comoving cubic box with periodic boundary conditions, following Ref. [81]. Spatial derivatives are computed using a Fourier pseudospectral method, and time integration is handled by a fourth-order Runge-Kutta scheme.
@@ 2.2 The initial conditions
$S(\xi,\vec{x})$ $=S_{i}(\vec{x}),$ (2.29)
$\Phi(\xi,\vec{x})$ $=\Phi_{i,\mathrm{ad}}(\vec{x})\left(\frac{8}{5\xi^{3}}(\sqrt{1+\xi}-1)-\frac{4}{5\xi^{2}}+\frac{1}{5\xi}+\frac{9}{10}\right)$
$+S_{i}(\vec{x})\left(\frac{16}{5\xi^{3}}(1-\sqrt{1+\xi})+\frac{8}{5\xi^{2}}-\frac{2}{5\xi}+\frac{1}{5}\right),$ (2.30)
@@
$S^{\prime}(\xi,\vec{x})$ $=0,$ (2.31)
$\Phi^{\prime}(\xi,\vec{x})$ $=f(\tau)\Phi_{i,\mathrm{ad}}(\vec{x})\left(\frac{8}{5\xi^{3}}-\frac{1}{5\xi^{2}}+\frac{4}{5\xi^{3}\sqrt{1+\xi}}-\frac{24(\sqrt{1+\xi}-1)}{5\xi^{4}}\right)$
$+f(\tau)S_{i}(\vec{x})\left(-\frac{16}{5\xi^{3}}+\frac{2}{5\xi^{2}}-\frac{8}{5\xi^{3}\sqrt{1+\xi}}-\frac{48(1-\sqrt{1+\xi})}{5\xi^{4}}\right),$ (2.32)
@@
The cosmological Gaussian random field $f_{g}$ is usually defined in momentum space [59]. Adopting the Fourier transfer convention
$f_{g}(\vec{x})=\int\frac{\mathrm{d}^{3}x}{(2\pi)^{3}}f_{g}(\vec{k})e^{i\vec{k}\cdot\vec{x}},$ (2.33)
and assuming the Fourier modes are complex Gaussian variables with a probability density function
$\mathscr{P}[f_{g}(\vec{k})]=\frac{1}{\pi\sigma_{k}^{2}}\exp\left\{-\frac{|f_{g}(\vec{k})|^{2}}{\sigma_{k}^{2}}\right\},$ (2.34)
leads to
$\langle f_{g}(\vec{k})f_{g}(\vec{q})\rangle=\sigma_{k}^{2}\delta^{3}(\vec{k}+\vec{q}),$ (2.35)
thus, we have $\sigma_{k}^{2}=(2\pi)^{3}\times(k^{3}/(2\pi^{2}))\mathcal{P}_{f_{g}}(k)$, where $\mathcal{P}_{f_{g}}(k)$ is the dimensionless power spectrum of $f_{g}(\vec{k})$. Therefore, after defining the dimensionless power spectrum of Gaussian random fields $\Phi_{g}$ and $S_{g}$ [^1] Generalizing to non-Gaussian fields is straightforward [81].
@@
one can obtain the initial value of $\Phi_{i,\mathrm{ad}}$ (“ad” means adiabatic) and $S_{i}$ [81], where the equation of state parameter $w$ is defined as $\omega=p_{(r)}/(\rho_{(r)}+\rho_{(m)})$ and $\zeta_{g}$ is the gauge invariant curvature perturbation. For the initial relative velocity $V_{\mathrm{rel}}$, combining Eqs. (2.25), (2.26), and the exact solutions (2.29) and (2.30) shows that during radiation-domination or matter-domination,
@@ 3.2 The peak structures of general initial conditions
The number of such peaks is governed by momentum conservation, as represented by the step function in Eq. (3.2), which requires
$|k_{i}-k_{j}|\leq k\leq k_{i}+k_{j}.$ (3.7)
Assuming $k_{j}=nk_{i}$ and $n>1$, the appearance of an additional peak requires $(1+n)/\sqrt{3}k_{i}\geq(n-1)k_{i}$, which yields the constraint $n\leq 2+\sqrt{3}$. However, in the isocurvature case, a suppression factor of the form $(k_{\mathrm{eq}}/k_{i})^{4}$ or $(k_{\mathrm{eq}}/k_{i})^{2}(k_{\mathrm{eq}}/k_{j})^{2}$ arises. This implies that certain peaks in the mixed initial condition exhibit an explicit dependence on $k_{\mathrm{eq}}$.
@@
Figure 3: The GW energy spectra with different initial conditions. All power spectra of initial curvature perturbations are composed of two peaks of width $\Delta=1/30$ located at $k_{*1}$ and $k_{*2}$. In the left panel, $k_{*2}=3k_{*1}$, while in the right panel, $k_{*2}=4k_{*1}$. “ad” and “iso” denote adiabatic and isocurvature perturbation, respectively.
To be specific, for the left panel of Fig. 2, we adopt $B_{1}=B_{2}=100$ in Eq. (3.4) to compute representative examples in isocurvature case; for the right panel, we choose $A=0.01$ in Eq. (3.2) for adiabatic perturbation and $B=100$ in Eq. (3.3) for isocurvature perturbation for numerical evaluation in mixed initial condition case. To investigate whether different initial conditions can be distinguished via the GW energy spectra, we plot in Fig. 3 the spectra induced by a curvature power spectrum featuring two peaks at $k_{*1}$ and $k_{*2}$. Imposing equal amplitudes for the first and last peaks, we find the amplitude of the additional intermediate peak to be largely unchanged. Thus, the initial conditions can be distinguished solely by the precise spectral shape around the peaks, a point also discussed in Ref. [25].
@@ 4.1 SIGW from eMD isocurvature
where $\gamma\sim 0.2$ in the radiation-dominated (RD) era, $H_{f}$ is the Hubble parameter at the time of PBH formation, and $\beta$ denotes the fraction of PBHs at formation. In the lattice simulation, we define and input a dimensionless quantity $\tilde{k}_{\mathrm{UV}}=k_{\mathrm{UV}}/k_{c}$, where $k_{c}$ denotes some characteristic scale. Assuming that during the early radiation-dominated era (eRD), the energy density of PBHs is negligible compared to radiation at the time of PBH formation, and that initial density fluctuations on scales larger than $k_{\mathrm{UV}}^{-1}$ can be neglected, which means $\delta\rho_{\mathrm{PBH},f}+\delta\rho_{r,f}=0$. [^2] Initial density fluctuations arise when adiabatic perturbations are considered. In contrast, isocurvature perturbations emerge following PBH formation, typically after the peak-like adiabatic perturbation has entered the horizon. Therefore, within linear perturbation theory, the adiabatic and isocurvature modes could be treated independently. Therefore, the initial isocurvature perturbation is given by
@@
Figure 4: Left: The GWs energy spectra sourced by PBH isocurvature perturbations. We have set $\tilde{k}_{\mathrm{UV}}=40$ in our simulations. Right: The energy density fraction $\Omega_{i}=\rho_{i=m,r}/(\rho_{m}+\rho_{r})$ of PBH and radiation. The dashed lines denote $\Omega_{m}$, while the solid line describes $\Omega_{r}$. $t$ is the cosmic time. The gray lines are (fitting) reference lines indicating specific power-law scalings. The gray line with $k^{11/3}$ and $k$ scalings are the analytical solutions derived in Ref. [24].
As illustrative examples, we consider PBH masses of $M_{\mathrm{PBH}}=10^{-4}g$ and $2\times 10^{-4}g$, and initial abundances $\beta=5\times 10^{-2},10^{-2},5\times 10^{-3}$ in our simulations [^4] For larger PBH masses or smaller initial abundances, the values of $t_{\mathrm{eva}}$ or $\tau_{*}$ in the simulation become very large. Simulating such cases is very challenging, which may require switching to a logarithmic time evolution scheme for the equations., as shown in Fig. 4. Our simulation has verified the $k$-scaling reported in Ref. [24]. However, while the near-peak power law in that work is $k^{11/3}$, our results indicate a possible dependence on the PBH decay time and the duration of the PBH-dominated era. It can be seen that a larger initial abundance or a larger PBH mass leads to a steeper power-law behavior near the peak, as illustrated in Fig. 5. An interesting finding is that the peak height (depicted in Fig. 5) is dictated solely by the PBH mass, while the location of the break in the power-law scaling is governed only by the initial abundance of PBHs. This behavior may also originate from the different decay times of PBHs and the duration of the PBH-dominated era, an interpretation that warrants further analytical and numerical investigation in future work.
Figure 5: Left: The GWs energy spectra sourced by PBH isocurvature perturbations for different PBH masses. Right: The GWs energy spectra sourced by PBH isocurvature perturbations for different PBH abundances. In both figures, we have multiplied a constant to make their ultraviolet part the same, and the gray lines are (fitting) reference lines indicating specific power-law scalings.
Figure 6: Left: The GWs energy spectra sourced by Q-balls isocurvature perturbations. We have set $\tilde{k}_{\mathrm{UV}}=40$ in our simulations. Right: The energy density fraction $\Omega_{i}=\rho_{i=m,r}/(\rho_{m}+\rho_{r})$ of Q-balls and radiation. The dashed lines denote $\Omega_{m}$, while the solid line describes $\Omega_{r}$. $t$ is the cosmic time. The initial abundance is set as $\beta=5\times 10^{-2}$. The gray lines are (fitting) reference lines indicating specific power-law scalings.
@@ 4.2 SIGW from general initial conditions
A full treatment of general initial conditions is considerably more complex; here, we provide only a brief extension of the PBH-dominated case. Strictly speaking, isocurvature perturbations only emerge subsequent to PBH formation, typically well after the peak of the adiabatic perturbation has entered the horizon. Thus, one ought to initialize the adiabatic and isocurvature fluctuations at distinct times. However, the solutions (2.29) and (2.30) are also attractors in the subhorizon regime, as noted in Ref. [53]. Therefore, minor variations in the initial conditions should have a negligible impact if PBHs form sufficiently before radiation-PBH equality; thus, as an illustrative example, we simply superimpose the initial adiabatic and isocurvature contributions as the initial condition. Assuming that PBHs are formed from a peak-like adiabatic perturbation centered at $k_{*}$, the ultraviolet cutoff scale $k_{\mathrm{UV}}$ can be related to $k_{*}$ via the identification $k_{*}=\mathcal{H}_{f}$
@@
For $\beta=10^{-2}$ and $\gamma=0.2$, one will obtain $k_{*}\simeq 2.7k_{\mathrm{UV}}$, which is shown in Fig. 7. As shown, the gravitational waves are initially generated by the adiabatic perturbation, as depicted by the blue line, and are subsequently sourced by the isocurvature perturbation. The rapid transition from the eMD to the RD era amplifies the GW signal, leading to a larger peak amplitude of the adiabatic part. This enhancement indicates that a general initial condition must be considered, rather than assuming a simple linear superposition of the adiabatic and isocurvature components. The middle peak corresponds to the peak structure discussed in Subsection 3.2. By using (3.6), one can find there will be three peaks located at $k_{UV},3.7/\sqrt{3}k_{UV},5.4/\sqrt{3}k_{UV}$.
Figure 7: Left: The GW energy spectrum sourced by the mixed initial condition at different time steps. The power spectrum of the initial adiabatic perturbation is the Gaussian bump power spectrum Eq. (3.2) with $A=0.01,\Delta=1/30$ and $\tilde{k}_{*}\equiv k_{*}/k_{c}=2.7\tilde{k}_{\mathrm{UV}}=54$, while the initial isocurvature power spectrum is Eq. (4.5) with $\tilde{k}_{\mathrm{UV}}=20$. $\tau_{i}$ is the initial comoving time. Right: The energy density fraction of PBH and radiation. The mass of PBH is $M_{\mathrm{PBH}}=2\times 10^{-4}g$ and the evaporation time is roughly $812\tau_{i}$.